A 1.77×10-1 g sample of HNO3 and a 5.21 g sample of NO react in a closed 8.91 L container at 470 K, according to the following balanced chemical equation:

2HNO3(g) + NO(g) → 3NO2(g) + H2O(l)

Calculate the PNO2 (in atm) in the container after the reaction has gone to completion.


1
Expert's answer
2022-01-17T23:52:02-0500

Moles of HNO3 = 1.77 × 10-1 / 63 =0.003 mol.


Moles of NO = 5.21 / 30 = 0.174 mol.


HNO3 is limiting reagent .


So, moles of NO2 formed = 3× 0.003 / 2 mol.

Moles of NO2 = 0.0045 mol.


From ideal gas equation

P = nRT / V = 0.0045 × 0.0821 × 470 / 8.91


P = 0.02 atm ( APPROX) Answer .